An ideal gas is enclosed in a container of volume V at pressure P. It is being pumped out of the container using a pump with stroke volume v. What is the final pressure in container after n-stroke of the pump(assume temperature remains same)1.P(V/v+V)^n2.PV/(V-v)^n3.P*(V/v)^n4.P(V/V-v)^n

Grade:11

1 Answers

Arun
6 years ago
Dear Aniket
Initial pressure = P   Volume = V Let the initial temperature be T
P V = n R T
stroke volume of the pump = v  (let v Assuming an isothermal process:
After the pumping out gas of volume v out of  V, the number of moles of the given gas left in the container = n *(V-v)/V
As Volume V of the container is constant,
P1 = P * (V-v)/V = P (1 - v/V)
Again when gas is pumped out pressure will become
P2 = P1 * (1 - v/V) = P (1- v/V)²
After n pumpings,  Pressure inside the container = P (1 - v/V)n
Regards
Arun (askIITians forum expert)